OpenStudy (anonymous):

If I have ((x+3)/(x+1)) - 2) / (x + 1). Is there a rule I can use to get rid of the fraction or x + 1?

OpenStudy (unklerhaukus):

use\[\frac {a+b}c=\frac ab+\frac bc\] \[\frac{\frac{x+3}{x+1} - 2 }{ x + 1}=\frac{\frac{x+3}{x+1}}{x+1}-\frac{2}{x+1}\]

OpenStudy (anonymous):

dang it thank you I was trying to do that the long way..

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (anonymous):

no prob, glad to help or at least too..